OpenStudy (anonymous):

how do i find a integer of a square root that has a decimal like 7.3

OpenStudy (anonymous):

So in other words without using a calculator you would like to know the answer to \[\sqrt{7.3}?\]

OpenStudy (anonymous):

yeah i dont know how to use calculatore properly and it needs to be the best estimate thats an integer to it

OpenStudy (anonymous):

its pre algebra i have to find the best integer estimate for each square

OpenStudy (anonymous):

so basicly i have to find the clostest integer to 7.3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If you're allowed to use a calculator I suggest learning how to use it. Otherwise it becomes very annoying to estimate a correct value. If you have absolutely no choice then here is how you do it: Lets make the question a bit easier as 7.3 is a hard number to find the exact square root of. Lets say we need to find \[\sqrt{7}\]. We know that \[\sqrt{4} = 2 \] and \[\sqrt{9} = 3\] meaning that your answer is somewhere in between 2 and 3. 7 is closer to 9, therefore we can assume that the is greater than 2.5. The problem now becomes how close to 2.5 is it. This is where a calculator comes in really handy cause there is no perfect way of getting the exact value. I'd estimate it to be something like 2.75.

OpenStudy (anonymous):

k thxs